OpenStudy (anonymous):

Jeanne wants to build a fence to enclose an area for a new rose garden. She can afford 150 feet of fencing. The length of the rectangular garden will be 5 feet more than the width, w. Which inequality best describes the possible width of her garden? A w + ( w + 5 ) ≤ 150 B 2w + 2(w + 5) ≥ 150 C 2w + 2(w + 5) ≤ 150 D w + ( w + 5 ) ≥ 150

OpenStudy (anonymous):

i think it will be 2w+2(w+5)≤150

OpenStudy (p0sitr0n):

w =width w +5 = length perimeter = 2 * (w + length), p =< 150 2 (w + (w+5)) =< 150 2w + 2(w+5) =< 150 c

OpenStudy (anonymous):

Thanks